lordtin:

A credit card had an APR of 11.91% all of last year, and compounded interest daily. What was the credit cards effective interest rate last year? A. 12.64% B.11.91% C.13.37% D.17.84%

LunaBechtold:

B bc it is the exact same number and didn't raise

Nightmoon:

To find the answer you'd need to use the formula of ~ (1+APR/n)^n-1 (Apr stands for Annual Percentage Rate) + the N = is the amount compounded interests.

toga:

To find the effective interest rate, we can use the formula: \( \text{Effective Annual Rate} = (1 + \frac{r}{n})^n - 1 \) Where: - r = annual nominal interest rate (APR) - n = number of compounding periods per year In this case, the annual nominal interest rate (APR) is 11.91% and it is compounded daily, so n = 365. Plugging in the values: \( \text{Effective Annual Rate} = (1 + \frac{0.1191}{365})^{365} - 1 \) After calculating, the effective annual rate comes out to be approximately 12.64%, so the answer is A. 12.64%.